OpenStudy (anonymous):

2log2 15-log2 5+log2 3

OpenStudy (anonymous):

Do you now the log properties?

OpenStudy (solomonzelman):

I will give you the rules you need: ~ \(\large\color{black}{ \log( a^{b} )=b\log(a)}\) ~ \(\large\color{black}{ \log( c )+\log(d)=\log(c \times d)}\) ~ \(\large\color{black}{ \log( c )-\log(d)=\log(c \div d)}\)

OpenStudy (anonymous):

Never mind Solomon told you the properties that pertain to this problem.